OpenStudy (anonymous):

what is the length of AC when the angle only says 6x and the side angle being BD is 5x + 3

OpenStudy (nubeer):

u need to put up the figure here.

OpenStudy (anonymous):

figure please

OpenStudy (anonymous):

@nubeer @luckythebest

jimthompson5910 (jim_thompson5910):

this is a parallelogram, so the opposite sides are congruent which means, AC = BD 6x = 5x + 3 solve for x and use it to find AC

OpenStudy (anonymous):

oh ok so opposite sides of a parallelogram are equal. Hence 6x = 5x + 3. find x; hence use AC = 6X.

OpenStudy (anonymous):

I'm confused and stuck on how to find the x with only 5x + 3 @luckythebest

OpenStudy (anonymous):

@potatoqueen he has given 6x also. The opposite sides of a parallelogram are equal and parallel. So actually AC = BD Hence 6x = 5x + 3. can u evaluate for x ?

OpenStudy (anonymous):

I'm not sure how to, I might be doing it wrong it's not 7x is it? or is it just 6x sorry btw @luckythebest

OpenStudy (anonymous):

its quite clear cut in the figure that AC is 6x.

jimthompson5910 (jim_thompson5910):

6x = 5x + 3 6x-5x = 5x + 3-5x .. Subtract 5x from both sides. what's next?

OpenStudy (anonymous):

so it's 18 @jim_thompson5910

OpenStudy (anonymous):

yes good answer :)

jimthompson5910 (jim_thompson5910):

x = 3 AC = 18 correct

OpenStudy (anonymous):

sorry can you help with this one @jim_thompson5910

OpenStudy (anonymous):

he has given that BDE similar to BCA, correct?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so u can say that CORRESPONDING SIDES ARE PROPORTIONAL. BD/BC = DE/CA = BE/BA. so BC = 125, AC = 150, BD = 4X+40 , ED = 8X+20. u can plug in the values and solve for X.

OpenStudy (anonymous):

I got 3.75 I did something wrong

OpenStudy (anonymous):

@luckythebest

OpenStudy (anonymous):

Lets see. BD/BC = DE/CA. 4X+40/125 = 8X+20/150 On cross multiplication - 150(4x+40) = 125(8x+20) OR 600x + 6000 = 1000x+2500 OR 3500 = 400x 35 = 4x. Hence X = 8.75.

OpenStudy (anonymous):

and this one how would i do this please help

OpenStudy (anonymous):

@luckythebest

jimthompson5910 (jim_thompson5910):

Hint:  the 3 triangles are similar, which means we can say p/h = h/q h^2 = pq